OpenStudy (anonymous):

how do you factor : 49+42a+9a^2

OpenStudy (lgbasallote):

(3a+7)^2

OpenStudy (anonymous):

do i have to do something different because of the was it's put?

OpenStudy (lgbasallote):

hmm you'll notice that the first term and the last term are both squares ...so your first guess is that it will be a perfect trinomial so you try that out first

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

you have to split the middle term [9a ^{2}+42a+49\] =9a^2+21a+21a+49 =3a(3a+7)+7(3a+7) =(3a+7)^2